Determine the nature of the solutions of:
4x^2-3x+6=0
Need help solving!

Answer by TimothyLamb(4379)

4x^2 - 3x + 6 = 0
---
the above quadratic equation is in standard form, with a=4, b=-3, and c=6
---
to solve the quadratic equation, plug this:
4 -3 6
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
Answer 1:
the vertex of the quadratic equation is a minimum point: ( 0.375, 5.4375 )
---
Answer 2:
the roots are complex:
root1 = 0.375 + (1.16592238)i
root2 = 0.375 - (1.16592238)i
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php